OpenStudy (anonymous):

i can't understand the Polar Coordinate Graph of this r=4+cos( a ) a: it is an angel

OpenStudy (anonymous):

What is it that you do not understand?

OpenStudy (anonymous):

how could i graph it??????

OpenStudy (anonymous):

Here goes an atempt with Excel Note that for each value of a you need to calculate the length of r

OpenStudy (anonymous):

i understand this part... but if we put each values of angels from 0 to 360 in the equation we will get the values of r like this r1= 6 , r2=2, r3=-2 ,r4=2, r5=6 and i can't under stand how he draw it like this

OpenStudy (anonymous):

But this not the same expression you posted first

OpenStudy (anonymous):

sorry but i have an exam i am so confused and i need help

OpenStudy (anonymous):

i mean if we must begin from 6 to 2 then we must draw it from 2 to -2 but why he draws it from 2 to 0 i mean do u know how could i draw inner loop i can't understand this inner loop if u know it can u explain to me how could i draw it from the begining to the end

OpenStudy (anonymous):

The graph you have is r=2+4cos(a). As far as cos(a) can take values between +1 and -1, there will be values of r<0 (which conceptually is wrong). In order to draw them what can be done is that when you have an angle that makes r<0 (and in this case that happens when cos(a)<-0.5) instead of drawing the distance from the origin following the angle, you have to follow the angle 180+a. For instance. At a=180 it happens that r=2+4cos(180)=-2. Just draw a line of length 2 from the origin following angle 180+180=360º=0º Another example:a=150º--->r=2-2sqrt(3)=-1.46--->draw a line of length +1.46 following angle 150+180=330º

OpenStudy (anonymous):

The rule can be summarized as follows: r>=0, draw r and use angle "a" r<0, draw |r| and use angle "a+180"

OpenStudy (anonymous):

thank u so much i can understand it now

OpenStudy (anonymous):

you are welcome and all the best with your exam